OpenStudy (dezandalf42714):

Please Help? Jackson put fifteen gallons of gasoline in his dad"s car and spent thirty dollars. Use a direct variation to determine how many gallons of gasoline he put in his mom"s car if he spent twenty-six dollars.

OpenStudy (xapproachesinfinity):

direct varation y=kx if y cost and x gallon units if he paid 30 dollar for 15 gallons then we have 30=k(15) k=30/15=2 now we know k y=2x <=== this direct varation we can use it to find if he spent 26 $ how many gallons? 26=2x solve for x

OpenStudy (xapproachesinfinity):

k=2 is the prince per unit of gasoline

OpenStudy (dezandalf42714):

is that 13 gallons?

OpenStudy (xapproachesinfinity):

price*

OpenStudy (xapproachesinfinity):

yes

OpenStudy (xapproachesinfinity):

correct

OpenStudy (dezandalf42714):

thanks!!!!!

OpenStudy (xapproachesinfinity):

np!